OpenStudy (anonymous):

How do you integrate 5e^0.1t ?

OpenStudy (tkhunny):

Try \(\int e^{t}\;dt\)

OpenStudy (anonymous):

Do we need to add C?

OpenStudy (tkhunny):

Always. Try \(\int 5e^{t}\;dt\)

OpenStudy (anonymous):

5e^t/5 + C?

OpenStudy (anonymous):

@tkhunny Come back! :P

OpenStudy (tkhunny):

No good. First, does 5e^t/5 mean \(5e^{t/5}\;or\;\dfrac{5e^{t}}{5} = e^{t}\). Go to more effort to write clearly. Second, \(\int 5e^{t}\;dt = 5\int e^{t}\;dt = 5e^{t} + C\) Okay, thy, \(\int e^{5t}\;dt\)

OpenStudy (anonymous):

I'm confused...

OpenStudy (tkhunny):

The constant out front does nothing. Just bring it along.

OpenStudy (anonymous):

So what is the answer? I'm totally lost!

OpenStudy (tkhunny):

Do you agree with these? \(\int e^{t}\;dt = e^{t} + C\) \(\int 5e^{t}\;dt = 5e^{t} + C\)

OpenStudy (anonymous):

Oh I get it now. Cheers!

OpenStudy (tkhunny):

We haven't done \(\int e^{5t}\;dt\).

OpenStudy (anonymous):

Could you just tell me the answer because I seriously don't get it? I'll make sure to do another question after this.

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Nope. You have to answer my questions.

OpenStudy (anonymous):

Because I thought it was 5 x 0.1e^0.1 + 1

OpenStudy (anonymous):

0.5e^1.1?

OpenStudy (anonymous):

Or is it 5/1.1 * e^0.1/1.1?

OpenStudy (tkhunny):

\(\int e^{t}\;dt = e^{t} + C\) \(\int 5e^{t}\;dt = 5e^{t} + C\) \(\int e^{5t}\;dt = \dfrac{1}{5}e^{5t} + C\) \(\int 5e^{5t}\;dt = 5\left(\dfrac{1}{5}\right)e^{5t} + C = e^{5t} + C\) Find the derivatives and prove it. Those are all the tools you need. \(\int 5e^{t/10}\;dt\) = ??

OpenStudy (anonymous):

e^t/10 + C?

OpenStudy (tkhunny):

Maybe. Are you just guessing?

OpenStudy (anonymous):

I think it is.

OpenStudy (anonymous):

Did something happen?

OpenStudy (tkhunny):

Are you just goading me into doing it? \(\int 5e^{t/10}\;dt = 5\int e^{t/10}\;dt = 5\left(\dfrac{e^{t/10}}{1/10}\right) + C = 50e^{t/10} + C\) Whoops! I guess you didn't prove it.

OpenStudy (anonymous):

Can I just come back when my mind is fresh? I'm not getting anything!

OpenStudy (tkhunny):

Try a different problem. This one has two moving parts. Get used to both individual moving parts and you'll begin to see it. Look long and hard at my four examples with all the 5's.